### v1

# Evaluating the Capability of Large-scale Language Models on Chinese Grammatical Error Correction Task

Abstract
Large-scale language models (LLMs) has shown remarkable capability in various of Natural Language Processing (NLP) tasks and attracted lots of attention recently. However, some studies indicated that large language models fail to achieve promising result beyond the state-of-the-art models in English grammatical error correction (GEC) tasks. In this report, we aim to explore the how large language models perform on Chinese grammatical error correction tasks and provide guidance for future work. We conduct experiments with 3 different LLMs of different model scale on 4 Chinese GEC dataset. Our experimental results indicate that the performances of LLMs on automatic evaluation metrics (e.g. $F_{\text{0.5}}$ score) falls short of the previous sota models because of the problem of over-correction. Furthermore, we also discover notable variations in the performance of LLMs when evaluated on different data distributions. Our findings demonstrates that further investigation is required for the application of LLMs on Chinese GEC task.
1 Introduction
Building on InstructGPTĀ Ouyang etĀ al. (2022), ChatGPT has demonstrated its powerful ability to understand complex instruction and generate reasonable responses on various of NLP tasks. Following the technical trajectory of ChatGPT, a significant number of high-quality LLMs have emerged in recent times in both academia and industry, such as LLaMAĀ Touvron etĀ al. (2023), ChatGLMĀ Du etĀ al. (2022) and PaLMĀ Anil etĀ al. (2023). Previous studies found that these LLMs have achieved great performance on a wide range of NLP tasks, including machine translationĀ Jiao etĀ al. (2023), named entity recognitionĀ Li etĀ al. (2023) and text summarizationĀ Yang etĀ al. (2023).
Certain studies have token comprehensive investigations into the performance of LLMs in the domain of English grammatical error correction, yielding some interesting findingsĀ Fang etĀ al. (2023); Wu etĀ al. (2023). LLMs are not able to outperform sota models in terms of automatic evaluation metrics. This is primarily because LLMs tend to make unnecessary modifications to make the input sentences more fluent, which may result in over correction problem, and in some cases, even alter the original semantics of the input sentences.
In this report, we aim to explore the performance of LLMs in Chinese GEC task. We conducted experiments on various LLMs to investigate the influence of model size on the GEC results. Additionally, we attempted different test dataset from various data sources to explore the impact of data distribution on the outcomes.
2 Experimental Setup
Dataset Data Source #Sents #Err. Sent. Avg. Length Avg. Edits Avg. Refs
NLPCC Chinese learner 2000 1983 29.7 2.1 1.1
MuCGEC Chinese learner 1136 1136 44.0 4.0 2.2
FCGEC Native speaker 2000 1101 55.8 1.0 1.3
NaCGEC Native speaker 500 482 56.2 1.2 1.2
Table 1: Dataset statistics.
2.1 Dataset
We conduct experiments on four Chinese GEC dataset to provide a comprehension demonstration of LLMs’ capability. The detailed statistics of these dataset are shown in TableĀ 1.
2.1.1 GEC data from Chinese learners
We apply the test set of NLPCC-2018Zhao etĀ al. (2018) and the validation set of MuCGECZhang etĀ al. (2022) for evaluation. These two dataset collect the grammar errors made by foreigners during their process of learning Chinese.
2.1.2 GEC data from Chinese native speaker examination
We apply the validation set of FCGECXu etĀ al. (2022) and the validation set of NaCGECMa etĀ al. (2022) for evaluation. These two dataset are collected from Chinese native speakers’ language examinations.
2.2 Model
We conduct experiments on 3 LLMs with different model scales:
    ChatGPT [^1] https://platform.openai.com/docs/api-reference: we evaluate the performance of ChatGPT with OpenAI’s official API. We choose gpt-3.5-turbo as the evaluated model, which stands out as the most advanced and specifically optimized for chat functionality.
    ChatGLM-6BĀ Du etĀ al. (2022): ChatGLM is an open bilingual language model based on GLM framework which is optimized for Chinese QA and dialogue and exhibits a robust capacity for Chinese understanding.
    LLaMA-7BĀ Touvron etĀ al. (2023): LLaMA is a collection of foundation LLMs ranging from 7B to 65B parameters proposed by Meta AI. we applied the 7B model for evaluation.
2.3 Evaluation Metric
We evaluate models’ performance with Precision, Recall and $F_{\text{0.5}}$ from word level and char level respectively.
We adopt the official implementation of MaxMatch ($M^{2}$)Ā Dahlmeier and Ng (2012) scorer to calculate word-level $F_{\text{0.5}}$ score and choose PKUNLP as our word segment tool. We apply ChERRANTĀ  [^2] https://github.com/HillZhang1999/MuCGEC/tree/main/ scorers/ChERRANT for char-level metric calculation.
2.4 Prompt
Considering the differences in performance of large language models, we designed different prompts for them. These prompts are roughly the same in semantics, but there are some differences in details. The prompts are shown in FigureĀ 1
Figure 1: The prompt for three LLMs.
2.5 Setting details
We set temperature to 0.6 when applying ChatGPT for a reliable generated result. For ChatGLM-6B and LLaMA-7B, we conduct experiments on 4 GeForce NVIDIA 3080 Ti GPUs.
3 Experiment results
NLPCC MuCGEC FCGEC NaCGEC
Word-level P R $F_{\text{0.5}}$ P R $F_{\text{0.5}}$ P R $F_{\text{0.5}}$ P R $F_{\text{0.5}}$
GPT-3.5-turbo 28.69 33.92 29.60 36.69 36.26 36.61 5.47 14.94 6.26 10.25 17.83 11.20
ChatGLM-6B 24.60 9.82 18.91 34.31 9.38 22.41 7.09 5.65 6.74 16.98 7.89 13.80
LLaMA-7B 17.63 7.98 14.2 24.54 10.02 19.03 3.42 3.00 3.33 8.44 4.75 7.30
Sota 50.56 25.24 42.11 - - - 57.00 34.89 50.59 54.88 42.38 51.82
NLPCC MuCGEC FCGEC NaCGEC
Char-level P R $F_{\text{0.5}}$ P R $F_{\text{0.5}}$ P R $F_{\text{0.5}}$ P R $F_{\text{0.5}}$
GPT-3.5-turbo 18.07 26.80 19.33 22.79 30.03 23.94 3.38 12.88 3.96 5.38 12.30 6.06
ChatGLM-6B 16.76 7.54 13.47 22.93 7.45 16.20 4.36 4.96 4.47 8.41 5.07 7.43
LLaMA-7B 8.90 5.43 7.89 13.37 7.04 11.34 1.43 1.78 1.49 2.83 2.17 2.67
Sota 50.09 32.09 45.04 54.69 29.10 46.51 55.94 35.58 50.20 52.83 40.43 49.78
Table 2: Experimental results.
The experiment results are shown in TableĀ 2. There are some results worthy of discussion.
First, different data sources result in distinct evaluation results. LLMs exhibit significantly superior performance when trained on Chinese learner data (NLPCC and MuCGEC), as opposed to Chinese native speaker examination data (FCGEC and NaCGEC). According to our observations, the grammatical errors made by Chinese learners primarily involve the misuse of similar words or phrases, rather than incorrect sentence structures. In contrast, GEC data from Chinese native speaker examination maintains a higher level of regularity and is consisted of more complex structural errors. It is noteworthy that there exists gaps between GEC data from Chinese examination and Chinese native speakers’ daily spoken habits.
Second, different model scales also lead to distinct performance. The unified trend is that ChatGPT performs similarly on Precision with other 2 smaller models while achieves significant improvement in Recall. This implies that the evaluated LLMs have similar error correction capability while their error detection ability differs a lot.
Third, there still exists great gaps between state-of-art models and LLMs on automated evaluation metrics. Previous workĀ Jiao etĀ al. (2023) has found the problem of over-correction for LLMs, which has also been noticed in our experiment.
What’s more, it is hard to explain why the char-level evaluation metrics is significantly lower than word-level evaluation metrics, which is not noticed in previous work.
4 Conclusion
In this report, we explore the performance of various LLMs on Chinese grammatical error correction task. Experimental results indicate that there still remains gap between LLMs’ performance and current sota models. Furthermore, the performance of different LLMs’ performance is greatly impacted by the distribution of test data. Future work can focus on addressing the over-correction problem of LLMs and explore the untapped potential of LLMs in the field of grammatical error correction tasks.